Sequences
Changing the subject of a formula
Probability
Factorising
Expanding Brackets
Brainteasers
100

Write the next three terms of the following sequence

7    10    13    16    19

22       25      28
100

Make d the subject of 

e= d + 5

d= e - 5

100

Leon throws a biased coin, the probability of getting tails is 0.4. 

Work out the probability of getting heads

0.6 

100

Factorise

3y + 15 

3(y + 5) 

100
Expand 


a(a - 9)

a2 - 9a

100

Can you find the 5th term in the following pattern

77,  49,  36,  18 

8 

200

A number sequence has nth term of 6n + 3 

Write down the first four terms of this sequence

9, 15, 21, 27

200

make u the subject of

v = u + 10t

u= v - 10t 

200
Matthew is playing darts. The probability that he hits a bullseye is 0.3. 


Matthew throws 2 darts. Find the probability he hits the bullseye with both darts.

0.3 x 0.3 = 0.09

200

Factorise 

18x + 9

9(2x + 1)

200

Expand

y(3y + 5)

3y2 + 5y

200

which two digits do you need to swap to make the solution to this maths problem correct?

571 + 962 = 840 

9 and 2 

571 + 269= 840

300
Find an expression for the nth term of this sequence


2      9      16      23      30

7n - 5

300

Given that x + y= 1 

what does y equal?

y= 1 - x 

300

A bag contains 600 coloured counters. The counters are yellow, brown or orange. 

There are 117 yellow counters in the bag. the probability that a brown counter is chosen from the bag is 0.35. 

calculate the number of orange counters in the bag

273


300

Factorise

27x - 45

9(3x - 5)

300

Expand 

9(2f - 3) 

18f - 27

300

Without reordering any letters, how many animals appear in the sequence below? 

PHEASANTORTOISEALION

6 

Pheasant, sealion, ant, seal, tortoise and lion 

400

Work out the nth term of the following sequence 

10    7    4    1    -2

-3n + 13 

400

make d the subject of 

c= 4d + 5

d= c - 5/4

400

Helen is taking part in a quiz on TV. The probability she answers a question correctly is 0.8. 

Helen is asked 2 questions, find the probability she answers both questions correctly.

0.8 x 0.8 = 0.64

400

Factorise fully

21x + 35y - 14z 

7(3x + 5y - 2z)

400
Expand and simplify 


3(x + 5) + 2(x + 1)

5x + 17

400

There is one single letter that you can use to replace a letter in each of the words below so the first one becomes a different food, the second one becomes a different body part and the third one becomes a different animal. 

What is the letter? 

- PEAS

- WAIST

- CAT

R 


(PEAR, WRIST, RAT)

500

Write an expression for the nth term of this sequence 

-25     -37      -49       -61

-12n - 13 

500

v = u + 10t

work out the value of v when u= 4 and t= 3

34

500

List all the outcomes when you roll a fair die and flip a coin at the same time. 

H1,  H2,  H3,   H4,   H5,   H6

T1,   T2,   T3,   T4,   T5,   T6

500

Factorise fully

8ap + 12cp - 4p2

4p(2a + 3c - p)

500

Expand and Simplify

4(x - 1) + 3(2x + 5) 

10x + 11

500

There are two commonly used two-letter words that share the same letters. If you take the opposites of those words you should now have two three-letter words. 

What are those new words?

YES and OFF.

the 2 letter words are NO and ON and their opposites are YES and OFF 

M
e
n
u